# DeepC4: Deep Conditional Census-Constrained Clustering for Large-scale Multitask Spatial Disaggregation of Urban Morphology

### v1

@@ 4.1 Comparing DeepC4, METEOR, and GEM
To facilitate a temporally consistent comparison in the succeeding discussion, we uniformly scaled up the outputs of the previous studies of METEOR and GEM using the multipliers, 1.37 and 1.05, to convert from 2012 and 2020, respectively, to 2022. These multipliers are based on the assumptions that the outputs of METEOR used the 2012 dwelling counts from the available census records during its time of analysis and that the outputs of GEM derived a 50-year linear trend for building counts. When applied, the corresponding converted outputs of METEOR and GEM were found to be comparable with the least possible percentage differences, confirming that these assumptions can be used reliably. We also compare the outputs of DeepC4 with those of METEOR and GEM in terms of their respective numerical estimates by urban morphology indicators and the quality of their spatial disaggregation, and discuss their respective accuracies, for cases wherein census records are available.
@@
When classified into settlement types, in Table 3, the numbers of rural and urban dwellings of the outputs of DeepC4 result in percentage errors of 0.87% and 1.8%, respectively, whereas those of GEM overestimate the rural dwelling by 8% and underestimate the urban dwelling by 13%. Similarly, both rural and urban occupants considered by the DeepC4 perfectly matches the official census records whereas those of GEM respectively deviated by 5.6% and -7%. Considering the number of buildings, the outputs of DeepC4 estimated $\sim$178k fewer rural buildings and $\sim$119k more urban buildings. This signifies the value of a more fine-grained census statistics at the level of sectors, effectively reducing the resulting uncertainty across resolutions, compared to GEM which used a single average household size for the urban and rural areas of the entire Rwanda and five rows of data on population counts at the provincial level (Paul et al., 2022). In addition, these findings also underscore the effect of the selection of the average household size at different levels of aggregation on the reliability and uncertainty of the resulting large-scale exposure and risk outputs, which have been shown to be important in characterizing the urban-rural divide and its relative regional degrees of physical vulnerability for any subsequent risk calculation.
@@ 4.1.1 By urban morphology indicators
In Table 4, while the total buildings across all general groups of macro-taxonomy classes observe minimal percentage differences of 1.8% to 5.1%, DeepC4 and GEM, on one hand, share a closer agreement for reinforced concrete, concrete block, brick masonry, and wood. On the other hand, DeepC4 and METEOR have more precise estimates for sundried blocks and rubble stones whereas, only under the categories of informal and wattle-and-daub, DeepC4 is less favorable when compared to GEM and METEOR. Despite the absence of building-level information to validate this distribution of macro-taxonomy classes, our results still highlight the observed sensitivity on the possible composition of macro-taxonomy classes of the building stock at a large scale using various approaches. Nevertheless, Table 4 show that outputs of DeepC4 and GEM have a more similar underlying expert belief system, except for the ’informal’ category, which can be because by the overlapping with other macro-taxonomy classes and the use of sector-level rural and urban household sizes, significantly affecting the allocation of the number of buildings under the said category in DeepC4.
@@
For the inferred distribution of building heights between DeepC4 and GEM, Table 5 shows that the more discrete height categories (i.e., H:1, H:2, and HBET:8+) achieve the lowest percentage differences up to 13%, while the categories of building heights with overlapping ranges (i.e., H:3, HBET:3-6, and HBET:4-7) have higher percentage differences up to 23%. When combined, the group of H:3, HBET:3-6, and HBET:4-7 corresponds to a total percentage difference of 31%, with DeepC4 estimating 6,930 more buildings. In addition, DeepC4 has higher estimates for taller buildings (i.e., H:2, H:3, HBET:3-6, and HBET:4-7) while GEM has only higher estimates for H:1. While DeepC4 incorporated the weak supervision from the available building-level height information from Bachofer et al. (2019), these results also reflect the uncertainty in associating lower and upper bounds of building heights with a specific number of floor levels, as shown in Table 2 in B. For example, H:1 may likely correspond to a building height of 3 meters, but, due to the uniqueness of the different wall types and macro-taxonomy classes, it can be possibly lower for the case of many informal settlements with weak structural integrity and higher for the case of more improved structural designs of buildings, such as reinforced concrete. Our results, however, provide an alternative distribution of building height classes that is both based on the expert belief system also used in Paul et al. (2022) and an extended benchmarking from the use of weak supervision from the available groundtruth data from Bachofer et al. (2019). Overall, we note that integrating expert belief systems (e.g., local understanding of the number of floor levels) and observational data (e.g., mapped building heights in meters) is a key challenge in characterizing the regional urban morphology for disaster risk assessment.
@@ 4.1.2 By spatial disaggregation quality
Figure 3: Comparison of the spatial disaggregation between DeepC4 and METEOR at 500-meter pixel. Calculated percentage differences (a) visualized on a map and (c) expressed as a histogram chart showing proportion of pixels per province. Positive percentage differences indicate that outputs of DeepC4 are larger than those of METEOR and negative otherwise. Categorical difference as (b) a map and (d) a histogram chart showing the proportion of pixels for each descriptive category.
@@
Although METEOR estimated $\sim$168k more buildings, Figure 3b and  3d show that outputs of DeepC4 occupy more 500-meter grid pixels by 32% (Kigali City) to 49% (Southern Province), which is indicative of the urban development captured by the EO data for the year 2022. While METEOR estimates a slightly higher number of buildings in 34% of all grid pixels in the Northern Province, the general trend across all provinces in DeepC4 still shows that 25% of all grid pixels in Eastern and Southern Provinces to 42% for Kigali City have more building counts than METEOR. It is important to note that Figure 3b and  3d illustrate only the differences on the space or grid pixels with which either DeepC4 or METEOR has disaggregated values, regardless of the magnitude of their differences. However, these results emphasize that the quality of large-scale exposure datasets also concerns not only the numerical accuracy of the disaggregated building counts but also the spatial accuracy as to where these buildings are located on the map, signifying the critical influence of urban dynamics and temporal consistency of EO data on their degrees of exposure to natural hazards.
@@
Furthermore, in terms of the magnitude of differences between DeepC4 and METEOR, Figure 3c shows that DeepC4 and METEOR result in uniformly distributed percentage differences. However, towards the left end or colored red in Figure 3c, it also indicates that the outputs of METEOR are significantly higher than those of DeepC4 for all surrounding and larger provinces except Kigali City, despite METEOR occupies less grid pixels than DeepC4 in the previous discussion of Figure 3d. Figure 3a also confirms that the province of Kigali City has more grid pixels that are colored blue or higher building counts from DeepC4, which is equivalent to the observed distribution of positive percentage differences in Figure 3c. These results imply that both the size or areal extent and the relative location (i.e., whether the region is located in the center or perimeter) affect the quality of spatial disaggregation. Particularly, the large areal extents of Nothern, Eastern, Western, and Southern Provinces provide more possible and uncertain locations for disaggregation, while the province of Kigali City with the lowest areal extent at the center has limited and constrained locations, resulting in the lowest categorical difference at 32% where an estimate of DeepC4 exists in Figure 3d.
@@
In Table 6, even at the province level, DeepC4 appears to have a set of more accurate numbers of dwellings for both urban and rural areas than those of GEM. For urban dwellings, GEM underestimates by 9% to 19% for all provinces except the South Province, which observes a large overestimation by 21%. For rural dwellings, GEM significantly overestimates Kigali City by 24%, followed by the East and West Provinces by 13% and 17%, respectively. Nevertheless, GEM achieves the most accurate numbers for rural dwellings for the North and South provinces, which are also the most precise estimates with the least percentage differences at 3.99% and 0.66%, respectively, when compared with the outputs of DeepC4. Our results show that constraining the spatial disaggregation by incorporating census statistics and encoding expert belief systems in a form of probabilistic relationships among various urban morphology indicators has effectively established a consensus while learning efficient reduced-dimension representations that are able to cluster under weak supervision of available building-level groundtruth information and infer the distribution of the characteristics of the urban layout.
@@
Figure 4: Learning curves. (a) reconstruction loss (i.e., how effective the autoencoder is in reducing the high-dimensionality into a meaningful reduced-dimension representations); (b) prediction loss (i.e., how accurate the resulting reduced-dimension representations are with respect to available building groundtruth info); (c-e) absolute proportion of true positives for each prediction task (roof, wall, and height); and (f-h) similar plots but accounting for the imbalanced number of target classes.
@@
Figure 5: Usefulness of partially available building-level groundtruth. (a-d) a set of prediction maps as a result of DeepC4; (e-g) corresponding precision maps; (h) reference background map with sector labels; (i-k) aggregated sector-level values of precision maps; and (l) reference location.
@@ 4.2 Assessing Overall Predictive Performance
To understand the reliability of the learning capability and predictive performance of our trained DeepC4 model, we present the learning curves in Figure 4 with an example set of prediction and precision maps in Figure 5 on the use of weak supervision of partially available building-level groundtruth of Bachofer et al. (2019).
@@
In Figure 4a and  4b, our trained DeepC4 model has converged to a set of stable reduced-dimension latent representations, resulting in a mean-squared error of 0.66 (a 44% reduction from the initially calculated reconstruction loss) and a Euclidean distance-based prediction loss of 3.376 (a 5.4% reduction from the initially calculated prediction loss). The observed fluctuations can be attributed to the varying sizes and non-uniform clustering behavior of sectors with different proportion of clusters. The reconstruction loss curve continues to decline minimally beyond the optimal epoch at 184 but the prediction loss curve starts to increase, which implies that an optimal reconstruction quality has been achieved already and any further iteration on the reconstruction would only result in a set of learned latent representations that are too sensitive to a deviating clustering behavior.
@@
Furthermore, our results in Figure 4c,  4d, and  4e show that, even at the initial clustering implementation, the incorporation of constraints has already enabled a desirable performance with respect to the cluster-level sizes, as indicated by the proportion of true positives respectively starting at 98.97%, 96.64%, and 95.34% and optimized at 99.03%, 96.45%, and 95.87% for roof, wall, and height. As the joint multitask learning continues, the trade-off exhibits an increasing trend in Figure 4c for roof, an increasing-to-decreasing trend in Figure 4d for wall, and an increasing-to-constant trend in Figure 4e for height.
@@
In the same way, evaluating the proportion of true positives while accounting for the imbalanced number of target classes resulted in a similar behavior, particularly for the wall and height, respectively in Figure 4g and  4h. However, Figure 4f shows that, in the earlier part of the joint multitask training, it achieves a more optimal performance and tends to become constant in the later part, which, in Figure 4c, still indicates an increasing absolute proportion of true positives.
@@
In addition, the magnitude of the weighted proportion of true positives in Figure 4g being higher than that of Figure 4c reflects the large difference in cluster sizes among roof classes. Similarly, Figure 4g and  4h having lower magnitudes than those of Figure 4d and  4e respectively suggests that the number of wall and height classes may be limited and overlapping, despite our efforts in linking different classifications between Bachofer et al. (2019) and the classes in Table 1. This implies that the optimality based on the imbalanced number of target classes and different standards for target classes, which is common in an urban morphology disaggregation task with overly dominating group of classes, is an important consideration in spatial disaggregation. Nevertheless, even though the partially available building-level groundtruth provides limited or weak supervision, the results of our 5-fold cross-validation implementation show a drop across all metrics by only 2%.
@@
To visualize these metrics on the map, Figure 5a to  5d show a set of prediction maps for the distribution of roof, wall, height, and macro-taxonomy classes over a few sectors in Kigali City, as labeled in Figure 5h and referenced in Figure 5l. Using the available building-level groundtruth information from which we trained our model using their calculated losses, Figure 5e to  5g shows that most grid pixels are accurately predicted (i.e., true positives in colored green) for maps of roof, wall, and height. Presenting these pixel maps as sector-level summaries in Figure 5i to  5k, our results also confirm that our trained DeepC4 model achieved a desirably improved performance with the clustering task for roof and height classes being notably more precise compared to that for wall.

### v2

@@ 4.1 Comparing DeepC4, METEOR, and GEM
Comprehensively comparing the outputs of DeepC4, METEOR, and GEM and disentangling the benefits of using DeepC4 and improved fine-grained inputs (e.g., average household size for urban and rural areas) are challenging and non-trivial because of the closed-source information on the methodological implementation of past studies and their original input data. Thus, in the succeeding discussions, the basis of our comparative analyses and discussions are the resulting outputs as compared to the available 2022 census-based aggregated information in Table 2, Table 3, and Table 6. In addition, because of the absence and inherent difficulty to maintain groundtruth information for other inferred outputs such as the distributions of general macro-taxonomy classes and height categories at the national level, we instead calculated pairwise percent differences to assess inter-model agreement among the outputs of DeepC4, METEOR, and GEM in Table 4 and Table 5.
Another key challenge is the temporal inconsistency among the outputs of DeepC4, METEOR, and GEM. Thus, we uniformly scaled up the outputs of METEOR and GEM using the multipliers, 1.37 and 1.05, to convert from 2012 and 2020, respectively, to 2022. These multipliers are based on the assumptions that the outputs of METEOR used the 2012 dwelling counts from the available census records during its time of analysis and that the outputs of GEM derived a 50-year linear trend for building counts. Although such a simple projection approach can further propagate uncertainties, the corresponding converted outputs of METEOR and GEM were found to be comparable, with the smallest possible percentage differences, confirming that these assumptions can be used reliably. We also compare the outputs of DeepC4 with those of METEOR and GEM in terms of their respective numerical estimates by urban morphology indicators and the quality of their spatial disaggregation, and discuss their respective accuracies, for cases wherein census records are available.
4.1.1 By building, dwelling, and occupant counts
@@
When classified into settlement types, in Table 3, the numbers of rural and urban dwellings of the outputs of DeepC4 result in percentage errors of 0.87% and 1.8%, respectively, whereas those of GEM overestimate the rural dwelling by 8% and underestimate the urban dwelling by 13%. Similarly, both rural and urban occupants considered by the DeepC4 perfectly matches the official census records whereas those of GEM respectively deviated by 5.6% and -7%. Considering the number of buildings, the outputs of DeepC4 estimated $\sim$178k fewer rural buildings and $\sim$119k more urban buildings.
This signifies the value of a more fine-grained census statistics at the level of sectors, effectively reducing the resulting uncertainty across resolutions, compared to GEM which used a single average household size for the urban and rural areas of the entire Rwanda and five rows of data on population counts at the provincial level (Paul et al., 2022). In addition, these findings also underscore the effect of the selection of the average household size at different levels of aggregation on the reliability and uncertainty of the resulting large-scale exposure and risk outputs, which have been shown to be important in characterizing the urban-rural divide and its relative regional degrees of physical vulnerability for any subsequent risk calculation.
@@ 4.1.2 By urban morphology indicators
In Table 4, while the total buildings across all general groups of macro-taxonomy classes observe minimal percentage differences of 1.8% to 5.1%, DeepC4 and GEM, on one hand, share a closer agreement for reinforced concrete, concrete block, brick masonry, and wood. On the other hand, DeepC4 and METEOR have more precise estimates for sundried blocks and rubble stones whereas, only under the categories of informal and wattle-and-daub, DeepC4 is less favorable when compared to GEM and METEOR.
Despite the absence of building-level information to validate this distribution of macro-taxonomy classes, our results still highlight the observed sensitivity on the possible composition of macro-taxonomy classes of the building stock at a large scale using various approaches. Nevertheless, Table 4 show that outputs of DeepC4 and GEM have a more similar underlying expert belief system, except for the ’informal’ category, which can be attributed to the overlapping with other macro-taxonomy classes and the use of sector-level rural and urban household sizes, significantly affecting the allocation of the number of buildings under the said category in DeepC4.
For the inferred distribution of building heights between DeepC4 and GEM, Table 5 shows that the more discrete height categories (i.e., H:1, H:2, and HBET:8+) achieve the lowest percentage differences up to 13%, while the categories of building heights with overlapping ranges (i.e., H:3, HBET:3-6, and HBET:4-7) have higher percentage differences up to 23%. When combined, the group of H:3, HBET:3-6, and HBET:4-7 corresponds to a total percentage difference of 31%, with DeepC4 estimating 6,930 more buildings. In addition, DeepC4 has higher estimates for taller buildings (i.e., H:2, H:3, HBET:3-6, and HBET:4-7) while GEM has only higher estimates for H:1.
@@
While DeepC4 incorporated the weak supervision from the available building-level height information from Bachofer et al. (2019), these results also reflect the uncertainty in associating lower and upper bounds of building heights with a specific number of floor levels, as shown in Table 2 in B. For example, H:1 may likely correspond to a building height of 3 meters, but, due to the uniqueness of the different wall types and macro-taxonomy classes, it can be possibly lower for the case of many informal settlements with weak structural integrity and higher for the case of more improved structural designs of buildings, such as reinforced concrete.
Our results, however, provide an alternative distribution of building height classes that is both based on the expert belief system also used in Paul et al. (2022) and an extended benchmarking from the use of weak supervision from the available groundtruth data from Bachofer et al. (2019). Overall, we note that integrating expert belief systems (e.g., local understanding of the number of floor levels) and observational data (e.g., mapped building heights in meters) is a key challenge in characterizing the regional urban morphology for disaster risk assessment.
@@ 4.1.3 By spatial disaggregation quality
Figure 3: Comparison of the spatial disaggregation between DeepC4 and METEOR at 500-meter pixel. Calculated percentage differences (a) visualized on a map and (c) expressed as a histogram chart showing the proportion of pixels per province. Positive percentage differences indicate that outputs of DeepC4 are larger than those of METEOR and negative otherwise. Categorical difference as (b) a map and (d) a histogram chart showing the proportion of pixels for each descriptive category. The color encoding of map and bar chart is shared between (a) and (c), and between (b) and (d).
@@
Although METEOR estimated $\sim$168k more buildings, Figure 3b and  3d show that outputs of DeepC4 occupy more 500-meter grid pixels by 32% (Kigali City) to 49% (Southern Province), which is indicative of the urban development captured by the EO data for the year 2022. While METEOR estimates a slightly higher number of buildings in 34% of all grid pixels in the Northern Province, the general trend across all provinces in DeepC4 still shows that 25% of all grid pixels in Eastern and Southern Provinces to 42% for Kigali City have more building counts than METEOR.
It is important to note that Figure 3b and  3d illustrate only the differences on the space or grid pixels with which either DeepC4 or METEOR has disaggregated values, regardless of the magnitude of their differences. However, these results emphasize that the quality of large-scale exposure datasets also concerns not only the numerical accuracy of the disaggregated building counts but also the spatial accuracy as to where these buildings are located on the map, signifying the critical influence of urban dynamics and temporal consistency of EO data on their degrees of exposure to natural hazards.
@@
Furthermore, in terms of the magnitude of differences between DeepC4 and METEOR, Figure 3c shows that DeepC4 and METEOR result in uniformly distributed percentage differences. However, towards the left end or colored red in Figure 3c, it also indicates that the outputs of METEOR are significantly higher than those of DeepC4 for all surrounding and larger provinces except Kigali City, despite METEOR occupies less grid pixels than DeepC4 in the previous discussion of Figure 3d. Figure 3a also confirms that the province of Kigali City has more grid pixels that are colored blue or higher building counts from DeepC4, which is equivalent to the observed distribution of positive percentage differences in Figure 3c.
These results imply that both the size or areal extent and the relative location (i.e., whether the region is located in the center or perimeter) affect the quality of spatial disaggregation. Particularly, the large areal extents of Nothern, Eastern, Western, and Southern Provinces provide more possible and uncertain locations for disaggregation, while the province of Kigali City with the lowest areal extent at the center has limited and constrained locations, resulting in the lowest categorical difference at 32% where an estimate of DeepC4 exists in Figure 3d.
@@
In Table 6, even at the province level, DeepC4 appears to have a set of more accurate numbers of dwellings for both urban and rural areas than those of GEM. For urban dwellings, GEM underestimates by 9% to 19% for all provinces except the South Province, which observes a large overestimation by 21%. For rural dwellings, GEM significantly overestimates Kigali City by 24%, followed by the East and West Provinces by 13% and 17%, respectively. Nevertheless, GEM achieves the most accurate numbers for rural dwellings for the North and South provinces, which are also the most precise estimates with the least percentage differences at 3.99% and 0.66%, respectively, when compared with the outputs of DeepC4.
Our results show that constraining the spatial disaggregation by incorporating census statistics and encoding expert belief systems in a form of probabilistic relationships among various urban morphology indicators has effectively established a consensus while learning efficient reduced-dimension representations that are able to cluster under weak supervision of available building-level groundtruth information and infer the distribution of the characteristics of the urban layout.
@@
Figure 4: Learning curves. (a) reconstruction loss (i.e., how effective the autoencoder is in reducing the high-dimensionality into a meaningful reduced-dimension representations); (b) prediction loss (i.e., how accurate the resulting reduced-dimension representations are with respect to available building groundtruth info); (c-e) absolute $p_{valid}$ for each prediction task (roof, wall, and height); and (f-h) similar plots but accounting for the imbalanced number of target classes.
@@
Figure 5: Usefulness of partially available building-level groundtruth. (a-d) a set of prediction maps as a result of DeepC4; (e-g) corresponding $p_{valid}$ maps; (h) reference background map with sector labels; (i-k) aggregated sector-level values of $p_{valid}$ maps; and (l) reference location.
@@ 4.2 Assessing Overall Predictive Performance
To understand the reliability of the learning capability and predictive performance of our trained DeepC4 model, we present the learning curves in Figure 4 with an example set of prediction and $p_{valid}$ maps in Figure 5 on the use of weak supervision of partially available building-level groundtruth of Bachofer et al. (2019).
In Figure 4a and  4b, our trained DeepC4 model has converged to a set of stable reduced-dimension latent representations, resulting in a mean-squared error of 0.66 (a 44% reduction from the initially calculated reconstruction loss) and a Euclidean distance-based prediction loss of 3.376 (a measurable 5.4% reduction from the initially calculated prediction loss). During the initial training phase, the reconstruction and clustering losses were of comparable orders of magnitude and converged to stable values. Consequently, the current unweighted formulation is sufficient to jointly achieve reconstruction quality and the learning of a meaningful reduced-dimensional latent representation within the proposed framework.
@@
In addition, the observed fluctuation trends of prediction loss in Figure 4b, which are more pronounced than those of the reconstruction loss in Figure 4a, can be attributed to the varying sizes and non-uniform clustering behavior of sectors with different proportions of clusters. In contrast, the smoother curve of reconstruction loss is inherent to the spatially continuous structure of EO satellite imagery signals, which vary gradually across neighboring pixels. The reconstruction loss curve continues to decline minimally beyond the optimal epoch at 184 but the prediction loss curve starts to increase, which implies that an optimal reconstruction quality has been achieved already and any further iteration on the reconstruction would only result in a set of learned latent representations that are too sensitive to a deviating clustering behavior.
@@
Furthermore, our results in Figure 4c,  4d, and  4e show that, even at the initial clustering implementation, the incorporation of constraints has already enabled a desirable performance with respect to the cluster-level sizes, as indicated by the $p_{valid}$ respectively starting at 98.97%, 96.64%, and 95.34% and optimized at 99.03%, 96.45%, and 95.87% for roof, wall, and height. As the joint multitask learning continues, the trade-off exhibits an increasing trend in Figure 4c for roof, an increasing-to-decreasing trend in Figure 4d for wall, and an increasing-to-constant trend in Figure 4e for height.
These reported high performances across all tasks are consistent with the observed high degree of class imbalance, in which a small number of classes dominate the majority of grid pixels, as previously indicated in Table 5, in addition to the influence of the mapping scheme of shared and overlapping characteristics between the building-level groundtruth information and desired indicators of urban morphology. Accordingly, this explains the use of a modified $p_{valid}$ considering such class imbalance, as emphasized in Equation 7, to more appropriately evaluate the overall predictive performance in Figure 4f,  4g, and  4h.
@@
In the same way, evaluating the $p_{valid}$ while accounting for the imbalanced number of target classes resulted in a similar behavior, particularly for the wall and height, respectively in Figure 4g and  4h. However, Figure 4f shows that, in the earlier part of the joint multitask training, it achieves a more optimal performance and tends to become constant in the later part, which, in Figure 4c, still indicates an increasing absolute $p_{valid}$.
@@
In addition, the magnitude of the weighted $p_{valid}$ in Figure 4g being higher than that of Figure 4c reflects the large difference in cluster sizes among roof classes. Similarly, Figure 4g and  4h having lower magnitudes than those of Figure 4d and  4e respectively suggests that the number of wall and height classes may be limited and overlapping, despite our efforts in linking different classifications between Bachofer et al. (2019) and the classes in Table 1. This implies that the optimality based on the imbalanced number of target classes and different standards for target classes, which is common in an urban morphology disaggregation task with overly dominating group of classes, is an important consideration in spatial disaggregation. Nevertheless, even though the partially available building-level groundtruth provides limited or weak supervision, the results of our 5-fold cross-validation implementation show a drop across all metrics by only 2%.
@@
To visualize these metrics on the map, Figure 5a to  5d show a set of prediction maps for the distribution of roof, wall, height, and macro-taxonomy classes over a few sectors in Kigali City, as labeled in Figure 5h and referenced in Figure 5l. Using the available building-level groundtruth information from which we trained our model using their calculated losses, Figure 5e to  5g shows that most grid pixels are accurately predicted (i.e., valid predictions in colored green) for maps of roof, wall, and height. Presenting these pixel maps as sector-level summaries in Figure 5i to  5k, our results also confirm that our trained DeepC4 model achieved a desirably improved performance with the clustering task for roof and height classes being notably more precise compared to that for wall.